# F-BIAS team

### v1

@@ Introduction
We have begun advertising our services cautiously, and it is likely that the current number of project requests discussed above does not yet fully represent the actual demand in the territory we aim to cover. As our resources are not matched to user requests, we envision that we will have to prioritize requests in a close future. Consequently, the criteria for accepting projects and establishing their priority order must be determined, and a selection committee composed of team members must be established.
Open-desk policy and organization tools
The organization of remote open-desks has been more challenging than expected. To not tax too much of the engineers time and still provide a quality service to the life scientists, the session schedule has to be optimized. Users not showing up, or data not transferred beforehand are the most common issues that waste time. The transfer of data before the session should be enforced for maximum efficiency of the remote session. The registration form could be more comprehensive to really prepare the session beforehand. Such a form could be built by taking inspiration from the questions that should be asked for a bioimage analysis project (35) but should take into account the fact that the user filling it might be quite inexperienced with quantitative approaches and may be unfamiliar with the vocabulary. Specializing a chatbot (36) to help the user filling the online form could be a solution to propose an interactive and user friendly registration and could even help to answer basic questions beforehand.
@@
As we grew, the need to share more than expertise arose. With the advent of deep learning, the availability of annotated training data has become a key factor for bioimage analysis. A distributed facility can capitalize on the diversity of dataset the analysts have access to, through the host institutes and the different equipment, samples and topics they have and focus on. In addition to sharing data, a distributed core facility will need to share code with its members, such as broadly applicable routines. The bioimage analysis community is used to distribute data and code publicly, but a distributed core facility offers the opportunity to share them early and privately with its members, while the code is being elaborated and the data is being annotated. A central resource for hosting code and data is therefore desirable for a distributed core. In our case, we primarily relied on the resources of FBI, particularly those from the FBI.Data project.
@@ Conclusion
F-BIAS started with the aim of federating analysts scattered across the territory and is now a working, distributed national core facility for Bioimage Analysis, that offers value to its users and members. Since its creation, F-BIAS has offered 35 consultations. With the new resources, we can target more than 10 collaborative projects before the end of 2024. The feedback we received from users through the consultations has been very positive and the collaborative projects we are working on are key to the research we contribute to. We are convinced that the success of this initiative is not particular to France and can be easily replicated in other conditions. Beyond the desirable key components listed in the previous section, we find the most important factor in the design of such an endeavor is to build and maintain a strong scientific added-value for its members. High quality bioimage analysis services fundamentally depend on analysts and their skills. It is therefore crucial to create a structure that can attract and retain analysts, and expand their knowledge and skills. Our commitment to maintaining scientific value for the analysts of F-BIAS prevents its erosion, sustaining the analysts’ interest in contributing to it since 2021. The services built upon F-BIAS benefit from the improving proficiency of the analysts, and in turn, the services are organized so that they foster the development of the analysts’ skills.
@@
How do the services that F-BIAS offers compare to the ones a user can find in a local core facility, if they have access to one? Based on our experience we observed the following. Consultations are more frequent in local cores compared to those of F-BIAS. This allows local users to return multiple times to their local open-desk sessions, refining the advice they receive as they implement it in their research. This effect of frequent open-desks suits the iterative nature of imaging-based projects better, and could also be implemented within F-BIAS with additional open-desk sessions, possibly with a rotation of the analysts. F-BIAS collaborative projects tend to be shorter in effort compared to local projects. This is due mainly to the splitting of one user request in several F-BIAS projects, built iteratively, with a go / no-go decision at the conclusion of each. For instance, ambitious questions start with a feasibility study that makes a first F-BIAS project, and is followed by another one for implementation if the first is successful. This reduces the risk for a user of paying a large amount of money upfront for an uncertain outcome. There are other kinds of services that can be found in local core facilities and that we do not offer yet. In addition to providing consultation or project service, a distributed facility is a resource for training both the analysts and the users. Most engineers are already giving regular individual or group training in their host institutes or via networks such as GloBIAS. Despite these efforts, there is still a huge demand for more training events from users (1, 11). A structure like F-BIAS could be utilized to collect funds and organize a few days of intensive image analysis classes, targeting locations where such courses are rare.
@@
Bioimage analysis services are perfectly compatible with a remotely operating core. They can be efficiently delivered thanks to modern IT tools, and the Covid crisis forced the integration of remote collaboration into our work culture. At the same time, they have a very strong value to research, as we could infer from the feedback of researchers that do not have access to them. Initiatives such as F-BIAS can easily leverage a few scattered resources to create significant value for the territory it supports simply by federating them. As a prospective catalyst for innovation and collaboration in bioimage analysis, we can envisage several extra activities beyond services to increase our value to the scientific community. F-BIAS is well-positioned to serve as an expert panel for industry partners, particularly those connected to FBI. Additionally, such infrastructures are well positioned to become a hub for additional resources dedicated to supporting the research of a country. They can become the recipient of infrastructure grants or resources from national research agencies. These resources will help bridge the gap in resources we observe with bioimage analysis while making them available nationwide.
@@ Acknowledgements.
F-BIAS is supported by the Agence Nationale de la Recherche through France-BioImaging (ANR-10-INBS-04) and the institutes that host its members, listed in the authors affiliations. MB is funded by LABEX Cortex (ANR-11-577 LABX-0042) of University Lyon 1, within the program “Investissements d’Avenir” (ANR-11-578 IDEX-0007). We are particularly grateful to the executive team of FBI, notably Caroline Thiriet, Jean Salaméro, Édouard Bertrand, Alexandre Philips and René-Marc Mège for their support and impulse in creating the structure. We thank our respective managers and board of directors that backed the creation of F-BIAS and encouraged us to contribute to it. We thank Laurent Essioux for critical comments on this manuscript.
@@
G.L. wrote the first draft of the article. J-Y.T. created the figures, reviewed and edited the article, integrating inputs and contributions from G.L., S.R., V.B., M.Amb., T.P., J.M., P.P-G. and A-S.M. All authors reviewed and validated the paper and contributed to the creation of the core and its activities.

### v2

@@ Introduction
We have begun advertising our services cautiously, and it is likely that the current number of project requests discussed above does not yet fully represent the actual demand in the territory we aim to cover. As our resources are insufficient for user requests, we envision that we will have to prioritize requests in the near future. Consequently, the criteria for accepting projects and establishing their priority order must be determined, and a selection committee composed of team members must be established.
Open desk policy and organization tools
The organization of remote open desks has been more challenging than expected. To not tax too much of the engineers time and still provide a quality service to the life scientists, the session schedule has to be optimized. Users not showing up, or data not transferred beforehand are the most common issues that waste time. The transfer of data before the session should be enforced for maximum efficiency of the remote session. The registration form could be more comprehensive to really prepare the session beforehand. Such a form could be built by taking inspiration from the questions that should be asked for a bioimage analysis project (36) but should take into account the fact that the user filling it might be quite inexperienced with quantitative approaches and may be unfamiliar with the vocabulary. Specializing a chatbot (37) to help the user filling the online form could be a solution to propose an interactive and user friendly registration and could even help to answer basic questions beforehand.
@@
As we grew, the need to share more than expertise arose. With the advent of deep learning, the availability of annotated training data has become a key factor for bioimage analysis. A distributed facility can capitalize on the diversity of dataset the analysts have access to, through the host institutes and the different equipment, samples and topics they have and focus on. In addition to sharing data, a distributed core facility will need to share code with its members, such as broadly applicable routines. The bioimage analysis community is used to distribute data and code publicly, but a distributed core facility offers the opportunity to share them early and privately with its members, while the code is being elaborated and the data is being annotated. A few users may require not to make anything related to their project public before it is published. A central resource for hosting code and data is therefore desirable for a distributed core. Currently, we use the public github repositories of the tools we work with. For collaborative work, we primarily rely on the resources of FBI, particularly those from the FBI.Data project, and RENATER (38) collaborative services (Table 3).
{tblr}
width = colspec = Q[252]Q[521]Q[25]Q[25], cells = m, fg=Tundora, rowodd = AthensGray, row1 = Nandor,fg=white, column1 = l, column2 = l, hlines, vlines = Black, vline1 = -LunarGreen,
Functionality & Tool deployed
Online chat service for instant communication A Mattermost instance, deployed in the Nantes FBI
Online meetings with F-BIAS members MS Teams private meeting with all members invited
User registration page for open desk session An interactive page on FBI website
Data transfer from user for an open desk session The FileSender tool of RENATER
Online F-BIAS open desk session Two online meeting rooms, using Rendez-Vous by RENATER
Documents keeping tracks of open desk session content Shared Google spreadsheet and documents
Table 3: The common tools currently used in F-BIAS for collaborative work and delivering services.
@@ Conclusion
F-BIAS started with the aim of federating analysts scattered across the territory and is now a working, distributed national core facility for Bioimage Analysis, that offers value to its users and members. Since its creation, F-BIAS has offered 35 consultations. With the new resources, we can work on about 5 to 10 collaborative projects per year. The feedback we received from users through the consultations has been very positive and the collaborative projects we are working on are key to the research we contribute to. We are convinced that the success of this initiative is not particular to France and can be easily replicated in other conditions. Beyond the desirable key components listed in the previous section, we find the most important factor in the design of such an endeavor is to build and maintain a strong scientific added value for its members. High quality bioimage analysis services fundamentally depend on analysts and their skills. It is therefore crucial to create a structure that can attract and retain analysts, and expand their knowledge and skills. Our commitment to maintaining scientific value for the analysts of F-BIAS prevents its erosion, sustaining the analysts’ interest in contributing to it since 2021. The services built upon F-BIAS benefit from the improving proficiency of the analysts, and in turn, the services are organized so that they foster the development of the analysts’ skills.
{tblr}
width = colspec = Q[100]Q[100]Q[100], cells = m, fg=Tundora, rowodd = AthensGray, row1 = Nandor,fg=white, column1 = l, column2 = r, column3 = r, hlines, vlines = Black, vline1 = -LunarGreen,
Collaborative projects & Local core facility F-BIAS
N. projects / year / analyst 8.2 ± 2.6 (N=6) 8.0 ± 0.0 (N=2)
Project effort 44.8 ± 43.3 hours (N=52) 48.8 ± 27.9 hours (N=5)
Project duration 33.9 ± 20.6 months (N=52) 7.3 ± 4.0 months (N=3)
Table 4: Statistics on collaborative projects comparing a local core facility (the Image Analysis Hub (IAH) of the Institut Pasteur) and F-BIAS. For the IAH data, we used the project management system deployed there (PPMS, Stratocore). These statistics are built on data from collaborative projects for local users and excludes other activities (consultations, maintaining computers and commercial software, administrative work, development projects, and meetings). To calculate the number of projects per analyst per year, we counted the active projects in 2024, including both ongoing and completed projects. We then determined the number of projects per analyst, weighted by the fraction of time each analyst dedicates to the IAH. The project effort and duration were measured for projects completed after 2022, from the number of hours billed to users. Duration was calculated as the time from project start to closure. For the F-BIAS column, we use the data from Table 2, using the billed time for the effort, and the actual duration of the 3 projects we completed. Two analysts dedicated 50% of their time to collaborative projects in 2024, which gives a number of projects per analyst of 8 with the 8 projects we worked on this year.
@@
How do the services that F-BIAS offers compare to the ones a user can find in a local core facility, if they have access to one? Based on our experience we observed the following. Consultations are more frequent in local cores compared to those of F-BIAS. This allows local users to return multiple times to their local open desk sessions, refining the advice they receive as they implement it in their research. This effect of frequent open desks suits the iterative nature of imaging-based projects better, and could also be implemented within F-BIAS with additional open desk sessions, possibly with a rotation of the analysts. F-BIAS collaborative projects are similar in effort compared to local projects, and national and local analysts will work on roughly the same amount of projects per year (Table 4). Local projects extend for much longer however, mainly because in the local core used in the comparison, projects stay open until they result in a publication when possible. Also, F-BIAS projects tend to be shorter in duration because user requests are divided into several iterative projects, each concluding with a go/no-go decision. For instance, ambitious questions start with a feasibility study that makes a first F-BIAS project, and is followed by another one for implementation if the first is successful. This reduces the risk for a user of paying a large amount of money upfront for an uncertain outcome.
There are other kinds of services that can be found in local core facilities and that we do not offer yet. In addition to providing consultation or project service, a distributed facility is a resource for training both the analysts and the users. Most engineers are already giving regular individual or group training in their host institutes or via networks such as GloBIAS. Despite these efforts, there is still a huge demand for more training events from users (1, 11). A structure like F-BIAS could be utilized to collect funds and organize a few days of intensive image analysis classes, targeting locations where such courses are rare. Local core facilities also offer access to shared workstations, with commercial and open-source software for image analysis installed and maintained. This facilitates autonomous image analysis by the users, and by pooling costly resources in the core facility, enables cost reduction for the host institute. A remote version of this service for open-source software would be highly valuable, considering the challenges of properly deploying these tools. For this a remotely operating core could leverage cloud-based solutions, deployed on computing resources paid for with a central budget, and maintained by a member of the national core with dedicated time for this task.
This prompts for a comparison between the national and local analyst roles. The services we offer in F-BIAS are currently limited to consultations and collaborative projects, while when we act as analysts for our local cores, we are also tasked with delivering the supplemental services listed above. But the main differences arise from daily interaction opportunities. A local analyst will have the chance to develop frequent formal and informal interactions with their users, in the scope of a collaborative project or in recurring consultations. This will help the analyst build deeper professional relationships with the local scientific community they support, and develop knowledge in the biological domain of application of the projects. This in turn can accelerate and enrich future collaborations with the same community. A national analyst engages with possibly many scientific communities (as can be inferred from Table 2). While this is stimulating, it poses a challenge in developing deep expertise across all these domains. Additionally, an analyst working remotely will interact with their users solely through the medium of the project and only for the duration of the project. This makes it difficult to build a sense of togetherness and shared purpose with the overarching research endeavor of an institute. These challenges are common to any team delivering remote services. In our current model, however, they have a minimal negative impact. F-BIAS is a virtual, distributed infrastructure and its members are part of research institutes from which they operate and reach a wide audience. As such we are also part of a local team that help mitigate the negative effects of remotely operating for analysts. As stated above this model comes also with straightforward project management, which is a significant benefit. Building on these advantages, local core facilities will continue hosting analysts funded by FBI, contingent on future developments. Another key connection for collaborative projects is with the local imaging and microscopy cores. Imaging-based projects are highly iterative, and the input of these experts is crucial for their success. It is important therefore that a user brings both analysts and imaging specialists in the project discussions, particularly when these discussions involve experimental design. Within the scope of F-BIAS, we have found no significant difference between local and remote projects on this topic. The efforts required to connect with imaging specialists or to have users involve them in discussions are similar for both local and remote projects.
F-BIAS will need to address several organizational points if its services are to be sustained over the long term. For instance, who will become responsible for maintaining a tool or pipeline we delivered when an F-BIAS analyst leaves the group? So far, this has not been required with the tools developed so far but this is an important aspect to consider for future projects. As stated above, the allocation of a project to an analyst is made considering availability and expertise, but what if a project with highly promising data and collaboration potential arises, and multiple experts are interested? These points overlap with those relevant to local core facilities. Since F-BIAS is a virtual infrastructure deeply rooted in these local cores, we will be able to draw inspiration from them to address these issues.
@@
Bioimage analysis services are perfectly compatible with a remotely operating core. They can be efficiently delivered thanks to modern IT tools, and the Covid crisis forced the integration of remote collaboration into our work culture. At the same time, they have a very strong value to research, as we could infer from the feedback of researchers that do not have access to them. Initiatives such as F-BIAS can easily leverage a few scattered resources to create significant value for the territory it supports simply by federating them. As a prospective catalyst for innovation and collaboration in bioimage analysis, we can envision several extra activities beyond services to increase our value to the scientific community. F-BIAS is well-positioned to serve as an expert panel for industry partners, particularly those connected to FBI. Additionally, such infrastructures are well positioned to become a hub for additional resources dedicated to supporting the research of a country. They can become the recipient of infrastructure grants or resources from national research agencies. These resources will help bridge the gap in resources we observe with bioimage analysis while making them available nationwide.
@@ Acknowledgements.
F-BIAS is supported by the Agence Nationale de la Recherche through France-BioImaging (ANR-24-INBS-0005 FBI BIOGEN, core funding, funds for M. Anselmet and A. Meslin) and the institutes that host its members, listed in the authors affiliations. M. Breuilly is funded by LABEX Cortex (ANR-11-577 LABX-0042) of University Lyon 1, within the program “Investissements d’Avenir” (ANR-11-578 IDEX-0007). We are particularly grateful to the executive team of FBI, notably Caroline Thiriet, Jean Salaméro, Édouard Bertrand, Alexandre Philips and René-Marc Mège for their support and impulse in creating the structure. We thank our respective managers and board of directors that backed the creation of F-BIAS and encouraged us to contribute to it. We thank Laurent Essioux for critical comments on this manuscript. We thank Nicolas Goudin for his involvement in the structure. The figures use icons that are derived from the work of Hilmy Abiyyu Asad (the Noun Project) and released under CC BY-3.0.
@@
G.L. wrote the first draft of the article. J-Y.T. created the figures, reviewed and edited the article, integrating inputs and contributions from G.L., S.R., V.B., M.Amb., T.P., J.M., P.P-G., A-S.M and M-S.P. All authors reviewed and validated the paper and contributed to the creation of the core and its activities.